OpenStudy (anonymous):

though(-5,0) perp.to y= -5/4 x= 5

OpenStudy (gorv):

slope of line perpendicular to this = 4/5 y=4x/5+c put (-5,0) here 0=4*(-5)/5+c 0=4*(-1)+c 0=-4+c c=4 y=4x/5+4